4. Monosodium glutamate (MSG), a food-flavor enhancer, has been blamed for “Chinese restaurant syndrome,” the symptoms of which are headaches and chest pains. MSG has the following composition by mass: 35.51 percent C, 4.77 percent H, 37.85 percent O, 8.29 percent N, and 13.60 percent Na. What is its molecular formula if its molar mass is about 169 g?
1
Expert's answer
2020-09-08T05:57:52-0400

For example, 100 g of MSG will contain 35.51/12.01 = 2.956 mol of C, 4.77/1.007 = 4.726 mol of H, 37.85/15.99 = 2.367 mol of O, 8.29/14.006 = 0.5918 mol of N, 13.60/22.98 = 0.5918 mol of Na.

Then all quantities will be divided by the smallest amount of moles.

"\frac{2.956}{0.5918} = 5\;C"

"\frac{4.726}{0.5918} = 8\; H"

"\frac{2.367}{0.5918} = 4\; O"

"\frac{0.5918}{0.5918} = 1\; N"

"\frac{0.5918}{0.5918} = 1\; Na"

So, the empirical formula is C5H8O4NNa.

The molar mass of the empirical formula will be calculated.

"(5\times 12.01) + (8 \times1.007) + (4 \times 15.99) + 14.006 + 22.98 = 169.05 \;g/mol"

The molecular formula of MSG is the same as the empirical formula C5H8O4NNa.

Answer: C5H8O4NNa